OpenStudy (anonymous):

How do you find the range/domain of a parabola? f(x)=2x^2-4x-5

OpenStudy (anonymous):

I know you do x=-b/2a so 4/4=1 as your vertex thingy

hero (hero):

To find the domain of any algebraic function, you assume that the domain is all real numbers while acknowledging any restrictions on the domain (if any). To find the range of a quadratic function, you consider the leading term. If the leading term is positive then the range will be \((-\frac{b}{2a}, \infty)\) If the leading term is negative, then the range will be \((-\frac{b}{2a}, -\infty)\)